OpenStudy (anonymous):

Which graph correctly solves the system of equations below? y = − x2 + 1 y = x2 − 4

OpenStudy (anonymous):

A)

OpenStudy (anonymous):

Is that x to the power 2, or 2 times x?

OpenStudy (anonymous):

\[y = x^2 + 1\]

OpenStudy (anonymous):

Right, got it.

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

D)

OpenStudy (anonymous):

It's definitely C. When you have a quadratic equation of the form ax^2 + bx + c = 0 where a, b and c are constants then a will be the degree to which the graph is stretched and if it is negative it will reflect the graph along the x-axis, b will shift the graph to the right if its sign is negative and vice versa and c is the y-intercept.

OpenStudy (anonymous):

Thankyou so much!

OpenStudy (anonymous):

This means the graph y = -x^2 + 1 Is an upside down parabola from the - sign Is not stretched (though this is not very relevant except in comparison) from the fact that a = 1 Cuts the y-axis at y = 1 Now both B and C have these basic features but looking at y = x^2 - 4 we see that this graph is an unstretched, right way up parabola with an intercept at y = -4 But B has a somewhat squashed (another way of stretching) while c isn't thus, it is c